OpenStudy (anonymous):

A cylindrical container with a cross-sectional area of 68.2cm2 holds a fluid of density 826kg/m3 . At the bottom of the container the pressure is 119kPa . Assume Pat = 101 kPa and h = 2.22m Find the pressure at the bottom of the container after an additional 2.55×10−3m3 of this fluid is added to the container. Assume that no fluid spills out of the container.

OpenStudy (anonymous):

I used the equation h=V/CSA to find out the height that was added to the fluid: \[h = \frac{ 2.55*10^{-3}m ^{3} }{ 0.682m ^{2} } = 0.0037m\] Then, I added it to the height and used proportionality: \[P = \frac{ \left( 2.22m+0.0037m \right)119kPa}{ 2.22m } = 119198Pa\] I was told it is wrong, but idk what im doing wrong.

OpenStudy (anonymous):

Looking at the question, the cylinder seems to be open from the top. Hence you must take into account atmospheric pressure when you write the total pressure at the bottom: \[P(at) + \rho gh = P(bottom) \] Hence, your proportionality equation is incorrect. Use the above equation to find the answer.

OpenStudy (anonymous):

I used the equation that you gave: \[P(bottom) = 101kPa + (826kg/m ^{3})(9.81m/s^{2})(2.2237m) = 281*10^{3} Pa\] The answer is wrong. It's supposed to be 1.22×105 Pa

OpenStudy (anonymous):

1.22*10^5 Pa

OpenStudy (anonymous):

Check your calculations. I am getting the right answer.